two point charges +8q and -2q are located at x=0 and x=l respectively.The location of the point on the x-axis ate which the net electric field due to these two point chares is zero-

Grade:12

1 Answers

Aastha
6 years ago
Let the point be at x distance from -2q charge. Then the distance between +8q and the point will be l+x.If the net electric field is zero then,E1 + E2 = 0k(8q)/(l+x)^2 = - k(-2q)/(x)^28q(x)^2 = 2q (l+x)^24(x)^2 = (l+x)^22x = l+xl=x